HUBBLE SPACE TELESCOPE DAILY REPORT #3074 PERIOD COVERED: 0000Z (UTC) 03/15/02 - 0000Z (UTC) 03/18/02 Daily Status Report as of 077/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of HST 8968 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} was used to characterize uncompensated environmental disturbances acting upon the HST during normal operation. The VDT is a passive test {not a forced-response test} used to obtain signatures for both externally induced {e.g. SA-3} and internally induced {e.g. NCC and ACS mechanisms} disturbances for comparison with past VDT results. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. The test occurs after release, and most of the VDT can be run during the BEA period. The -V1 sunpoint portion of the VDT occurs after the BEA period is complete. There were no reported problems. 1.2 Completed STIS/MA1/MA2 8959 (MAMA Dark vs. Temperature) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to measure the dark rate for both MAMA detectors as a function of instantaneous temperature and recent thermal history. These will be used to make measurements at the beginning and end of three consecutive SAA-free orbit blocks. The proposal completed with no reported problems. 1.3 Completed Six Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.4 Completed Thirteen Sets of ASC/HRC/WFC 9005 (CCD Functional) The Advanced Camera for Surveys (HRC and WFC) was used to measure the baseline performance and commandability of the CCD subsystems using exposures of bias frames, darks and flat fields with tungsten lamp. The on-board baseline compression algorithm for WFC data are measured. This procedure is carried out purely with internal targets and observations do not need to be contiguous. No anomalous activity occurred. 1.5 Completed Seven Sets of FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1 was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.6 Completed Five Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.7 Completed STIS/CCD/MA1/MA2 8960 (Jitter Test) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to image a point source with the FUV MAMA detector in TIMETAG mode, using power spectra of image position versus time to look for jitter at frequencies from about 0.2 to 100 Hz. Local count rate must be below the normal screening limit of 100 counts per second per pixel. Global count rate from the source will be roughly 400 counts per second over the entire detector. The global count rate from geocoronal Lyman-alpha will be maximized by observing a relatively bright source through a low transmission filter. The jitter before and after NCC is turned on will be compared. The proposal completed with no reported errors. 1.8 Completed Six Sets of NCS 8967 (Start NICMOS Cooling System) The NICMOS Cooling System was configured to re-cool the NICMOS detectors, to determine the optimal science operating temperature for the NICMOS detectors, and to verify stable operation. The goals are as follows: 1} cool the NICMOS to the coldest possible temperature as specified by the NICMOS team {73K} with HST at a hot attitude and the NCC compressor operating near its maximum allowable speed. 2} measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 3} demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature and 4} demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. All tests were successful. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 24 Successful: 24 Scheduled Re-acquisitions: 3 Successful: 3 2.2 FHST Updates: Scheduled: 36 Successful: 36 2.3 Operations Notes: HST operations were conducted from the GSFC STOCC throughout the weekend, and will remain there until at leat Wednesday. Using ROP SR-1A, the SSR-1 EDAC error counter was cleared once. The SSR-3 EDAC error counter was cleared twice per ROP SR-9A. The NICMOS Main Electronics Board, Side 2 (NMEBTMP2) flagged out-of-limits high at 074/190318Z with a value of 30.0645 degC (the limit is 30.0 degC). HSTAR 8553 was written. The ACS Cal Lamp -5V (JCLM5V) went out-of-limits high at 074/195613Z with a reading of -4.84332 volts. HSTAR 8556 documented the event. SSA transmitter was turned on and of for the intervals 074/2209Z to 074/2238Z, 074/2353Z to 075/0021Z, 076/0328Z to 076/0333Z and 076/0931Z to 076/0947Z as directed by ROP IC-2. HSTAR 8557 was produced when the FHST #2 mounting bracket temperature (TFHST2MB) went out-of-limits low with a value of -3.9 decC (limit is -5 degC) at 075/0000Z. The primary heater were known to have failed and the secondary heaters turned on at -2.9 degC. Operations requests were used to command the following: NSSC-1 HV Protect Table Configuration @ 075/0307z HV Protect Enable @ 075/0310z Power Off FHSTs for NCS activities @ 075/0454z NCC Circulator Loop Fill @ 075/0756z NCS CPL Standby to Pressure Prime @ 075/0758z NCS Pressure Prime to PCE On @ 075/1202z Temporary FHST Temperature Limits Off @ 075/1552z EPS Reconfiguration for NCS Turn On @ 075/2002z CS SDH Test @ 075/2142z NCC Circulator Start @ 075/2319z CS SDH TEST2 @ 075/2337z (collection on CCS-A) CS SDH TEST2 @ 075/2345z (collection on CCS-A/B/C) NCC Compressor Start @ 076/0015z NCS CPL to Control @ 076/0120z 8051 History Buffer Dump @ 076/0300z NCS Control to PID Operate @ 076/0339z Turn on NCS CPL Startup heater @ 076/0855z Turn on PCE @ 076/0945z Test CS SDH data collection @ 076/1307z, 076/1545z, 076/1633z, and 076/1711z The NSSC-1 status buffer was dumped at 075/2218Z, using ROP NS-03. Vehicle Disturbance Test #3 (Part 1) executed nominally between 076/045000Z and 076/095000Z. This portion of VDT #3 was a contingency test designed to be executed in real-time to capture NCC surging disturbances immediately after the transition to NCC PID Operate, at the maximum compressor speed during surging. This test was a 3-orbit test at +V3 sunpoint. Slightly more than one orbit of surging data was collected by the time the NCC safing event occurred at 076/072522Z. Analysis of the surging data revealed similar surging disturbance characteristics observed during the Vibration Emittance ground test performed in July 2000. Analysis of this flight data is ongoing. Vehicle Disturbance Test #3 (Part 2) executed nominally between 076/130500Z and 076/185000Z. This portion of VDT #3 was intended to acquire HST dynamic responses due to disturbances at +V3 sunpoint during NCC surging at the maximum compressor speed. Due to the NCC safing, the compressor has been off. Therefore, the data acquired during this VDT was similar in character to data acquired during VDT #2. Per HSTAR 8564, the ACS input voltage (J28VIN1) broke its lower limit at 076/035743Z and continued to break it intermittently until 076/071145Z. The limit is 24.5 volts. Low battery and spacecraft bus parameters were noticed at eclipse end between 076/053104Z and 076/053508Z. HSTAR 8573 was written. The NICMOS Cryo Cooler (NCC) safed (at 076/0725Z) during a ZOE. At AOS (076/074734Z) four new status buffer messages were observed. HSTAR 8565 was generated. The NCC remains in this state pending furhter analysis. At 076/085821Z, the battery #5 current (CBAT5CUR) went out-of-limits and remained in that state until 076/090108Z. HSTAR 8566 was written. Per ROP DF-09A, t-he telemetry format was changed via real-time commanding at 076/1524Z and at 076/1852Z. NCS CPL flooding started during start-up at 076/2200Z. It was observed during the initial cold shock that the NCS CPL lop flooded, not an unexpected state. CPL recovered without ground intervention. No NCS Commanding was impacted. HSTAR 8561 was generated. There was an HST coarse attitude error of 2.31 degrees as of 077/0825Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW